A block of copper of unknown mass has an initial temperature of 62.4 ∘C. The copper is immersed in a beaker containing 92.1 g of water at 22.6 ∘C. When the two substances reach thermal equilibrium, the final temperature is 25.0 ∘C. What is the mass of the copper block? Assume that Cs,H2O=4.18J/(g⋅∘C) and Cs,Cu=0.385J/(g⋅∘C).

Express the mass to two significant figures and include the appropriate units.
Answer #1

m(water) = 92.1 g

T(water) = 22.6 oC

C(water) = 4.18 J/goC

m(copper) = to be calculated

C(copper) = 0.385 J/goC

T = 25.0 oC

We will be using heat conservation equation

use:

heat lost by copper = heat gained by water

m(copper)*C(copper)*(T(copper)-T) = m(water)*C(water)*(T-T(water))

m(copper)*0.385*(62.4-25.0) = 92.1*4.18*(25.0-22.6)

m(copper)*14.399 = 923.9472

m(copper)= 64.2 g

Answer: 64 g